Calculate the molar mass of the following compounds. Show your work.


Water, H2O 

carbon dioxide, CO2 

sodium chloride, NaCl 

calcium hydroxide, Ca(OH)2 

potassium permanganate, KMnO4




1
Expert's answer
2021-02-22T05:53:00-0500

Solution:

M - molar mass

Ar - standard atomic weight


1) Water, H2O:

M(H2O) = 2×Ar(H) + Ar(O) = 2×1.00784 + 15.999 = 18.01468 = 18.015

M(H2O) = 18.015 g/mol


2) Carbon dioxide, CO2:

M(CO2) = Ar(C) + 2×Ar(O) = 12.0107 + 2×15.999 = 44.0087 = 44.009

M(CO2) = 44.009 g/mol


3) Sodium chloride, NaCl:

M(NaCl) = Ar(Na) + Ar(Cl) = 22.9898 + 35.453 = 58.4428 = 58.443

M(NaCl) = 58.443 g/mol


4) Calcium hydroxide, Ca(OH)2:

Ca(OH)2 = CaO2H2

M(Ca(OH)2) = Ar(Ca) + 2×Ar(O) + 2×Ar(H) = 40.078 + 2×15.999 + 2×1.00784 = 74.092

M(Ca(OH)2) = 74.092 g/mol


5) Potassium permanganate, KMnO4:

M(KMnO4) = Ar(K) + Ar(Mn) + 4×Ar(O) = 39.0983 + 54.938 + 4×15.999 = 158.032

M(KMnO4) = 158.032 g/mol